### v2

# Exponential Sums with Sparse Polynomials and Distribution of the Power Generator

Abstract.
We obtain new bounds on complete rational exponential sums with sparse polynomials modulo a prime, under some mild conditions on the degrees of the monomials of such polynomials. These bounds, when they apply, give explicit versions of a result of J. Bourgain (2005). In turn, as an application, we also obtain an explicit version of a result of J. Bourgain (2010) on rational exponential sums with sparse polynomials modulo an arbitrary composite number. We then use one of these bounds to study the multidimensional distribution of the classical power generator of pseudorandom numbers, which has not been possible within previously known results.
Key words and phrases:
Complete rational sums, sparse polynomials, power generator, pseudorandom numbers
2020 Mathematics Subject Classification:
11L07, 11K45, 11T23, 65C10
1. Introduction
1.1. Set-up and motivation
Let $q$ be a positive integer. Given a polynomial $f\in{\mathbb{Z}}_{q}[X]$ over the residue ring we define the following exponential sums
(1.1) $S_{q}(f)=\sum_{x\in{\mathbb{Z}}_{q}^{*}}{\mathbf{\,e}}_{q}(f(x)),$
over the group of units ${\mathbb{Z}}_{q}^{*}$, where
${\mathbf{\,e}}_{q}(z)=\exp(2\pi iz/q),$
and we always assume that the elements of ${\mathbb{Z}}_{q}$ are represented by the set $\{0,\ldots,q-1\}$.
Here we are mostly interested in sparse polynomials, that is, polynomials of the form
(1.2) $f(X)=\sum_{i=1}^{r}a_{i}X^{e_{i}}\in{\mathbb{Z}}_{q}[X],$
with a small (compared to the degree) number $r$ of distinct exponents $e_{r}>\ldots>e_{1}\geq 1$, for which we always assume
(1.3) $\gcd(a_{1},\ldots,a_{r},q)=1$
We first discuss the case of prime $q=p$, where we use ${\mathbb{F}}_{p}={\mathbb{Z}}_{p}$ to denote the finite field of $p$ elements.
The celebrated result of Weil [43] asserts that for any non-constant polynomial $f\in{\mathbb{F}}_{p}[X]$, over a finite field ${\mathbb{F}}_{p}={\mathbb{Z}}_{p}$ of $p$ elements, where $p$ is prime, we have
(1.4) $\left|S_{p}(f)\right|\leq(d-1)p^{1/2},$
where $d=\deg f$, see also, for example, [21, Chapter 11] and [25, Chapter 6].
Clearly the bound (1.4) becomes trivial for $d\geq p^{1/2}+1$, which is the limit of our present knowledge. However, in the case of sparse polynomials (1.2) where we can also assume $p-1\geq e_{r}>\ldots>e_{1}\geq 1$ and $a_{1},\ldots,a_{r}\in{\mathbb{F}}_{p}^{*}$, better bounds are known.
Most of attention was dedicated to the monomial case, that is, to $r=1$. Since the first improvement of the general bound (1.4) given in [35], several stronger bounds have been obtained, see, for example [20, 23, 33] and references therein. We observe, that if $r=1$, then without loss of generality, we can assume that $d=e_{1}\mid p-1$, which reduces the corresponding sums to estimating exponential sums of elements of subgroups of ${\mathbb{F}}_{p}^{*}$ of order $(p-1)/e$. In particular, by a result of Shkredov [33, Theorem 1], for $f(X)=aX^{d}$ with $a\in{\mathbb{F}}_{p}^{*}$, we have
(1.5) $S_{p}(f)\ll d^{1/2}p^{2/3}(\log p)^{1/6},$
where, as usual, the notations $U=O(V)$, $U\ll V$ and $V\gg U$ are equivalent to $|U|\leq cV$ for some positive constant $c$, which throughout this work may depend only on $r$ (and thus the implied constant in (1.5) is absolute). Note that in some ranges, stronger bounds are available, but (1.5) is quite sufficient for our purpose.
For $r=2$ a large variety of bounds is available [19, 28, 29, 37, 38], which require rather strong restrictions on some greatest common divisors associated with the powers $d$ and $e$ of monomials in $f(X)=aX^{d}+bX^{e}\in{\mathbb{F}}_{p}[X]$. For our purpose we use a result of Cochrane and Pinner [9, Theorem 1.3]:
(1.6) $|S_{p}(f)|\leq\gcd(d-e,p-1)+2.292g^{13/46}p^{89/92}$
where $g=\gcd(d,e,p-1)$.
For larger values of $r$ there are also some estimates that go beyond the Weil bound (1.4), see [8, 26, 38]. However, they apply only to rather special choices of the exponents $e_{1},\ldots,e_{r}$.
The case of arbitrary polynomials $f$ as in (1.2), under some natural condition (which is essentially necessary as well) has been treated by Bourgain [3]. Namely, for arbitrary $r\geq 2$ and arbitrary polynomials $f$ as in (1.2), under a natural condition that
$\gcd(e_{i},p-1),\,\gcd(e_{i}-e_{j},p-1)\leq p^{1-\varepsilon},\qquad i,j=1,\ldots,r,\ i\neq j,$
for some fixed $\varepsilon>0$, Bourgain [3, Theorem 1], using methods of additive combinatorics, has obtained the bound
(1.7) $\left|S_{p}(f)\right|\leq p^{1-\delta},$
with some $\delta>0$ depending only on $r$ and $\varepsilon>0$. Unfortunately, Bourgain [3] has not provided any explicit formulas for $\delta$ as a function of $r$ and $\varepsilon>0$, while an attempt to do so in [30, Theorems 4 and 5] seems to be faulty. In particular, it appears that the argument in [30] misquotes the result of [34, Corollary 16], which, after correcting, leads to exponentially smaller saving (and yet the rest of [30] is still to be verified).
For composite $q$ and “dense” polynomials $f(X)=a_{d}X^{d}+\ldots+a_{1}X\in{\mathbb{Z}}_{q}[X]$ of degree $d$ with $\gcd(a_{1},\ldots,a_{r},q)=1$, we have
$S_{q}(f)\ll q^{1-1/d},$
see, for example, [10, 39]. Note that this is a slight improvement of the classical bound of Hua (containing an additional factor $q^{o(1)}$), see, for example, [42, Theorem 7.1]. For sparse polynomials as in (1.2) and (1.3), but of fixed degree $\deg f=e_{r}$ a better bound of the type
(1.8) $|S_{q}(f)|\leq q^{1-1/r+o(1)}$
is given in [36, Theorem 1].
Bourgain [4], using the bound (1.7), has given several estimates for the sums $S_{q}(f)$ with sparse polynomials as in (1.2) and (1.3), For example, by [4, Theorem 1], for any $\eta>0$ there exist $\delta>0$ depending only on $r$ and $\eta>0$, such that for
(1.9) $\left|S_{q}(f)\right|\leq q^{1-\delta},$
provided that
$q\geq\exp\left(d^{\eta}\right),$
where $d=\deg f=e_{r}$. In particular, the derivation of (1.9) in [4] is based on (1.7) from [4]. Here, despite that our explicit version of (1.7) is more restrictive in terms of the conditions on the exponents $e_{1},\ldots,e_{r}$, it is sufficient to get a version of (1.9) with an explicit value of $\delta$. A further improvement comes from the new way we treat exponential sums modulo $p^{m}$, when $m\geq 2$ is not too large, see Lemma 4.2 below.
1.2. Applications
We also recall that the bound (1.6) has been used in [27] to study the uniformity of distribution properties of the power generator of pseudorandom numbers, More precisely, given an integer $e\geq 2$ and a prime $p$, this generator outputs the sequence $\{u_{n}\}$, defined by
(1.10) $u_{n}\equiv u_{n-1}^{e}\pmod{p},\quad 0\leq u_{n}\leq p-1,\qquad n=1,2,\ldots\,,$
with the initial value $u_{0}\not\equiv 0\pmod{p}$. Since the power generator has extensive applications in cryptography, and thus has been investigated in a number of works, see, for example, [3, 7, 12, 13, 15, 17, 16, 23, 24, 32] and the references therein, we also refer to surveys [41, 44].
We use our new bounds on exponential sums with sparse polynomials to study simultaneous distribution of $s$-tuples of the power generator, while the previous results allow explicit bounds only in the one-dimensional case as in [27]. We also discuss the introduced in [27] multivariate generalisations of the power generator to which our approach applies as well.
2. Main results
2.1. Exponential sums with sparse polynomials over ${\mathbb{F}}_{p}$
Here we use the bounds (1.5) and (1.6) as the basis for the inductive argument of Bourgain [3], in the form outlined by Garaev [18, Section 4.4] to derive a bound on exponential sums $S_{p}(f)$ with sparse polynomials $f$ as in (1.2).
For a real $\varepsilon>0$, we set
(2.1) $\kappa_{1}(\varepsilon)=\kappa_{2}(\varepsilon)=\varepsilon,$
and define the sequence $\kappa_{r}(\varepsilon)$, $r=3,4,\ldots$, recursively as follows
(2.2) $\kappa_{r}(\varepsilon)=\frac{\varepsilon}{t_{r}(\varepsilon)},$
where
(2.3) $t_{r}(\varepsilon)=\left\lceil\frac{r-2-\varepsilon}{2\kappa_{r-1}(\varepsilon)}\right\rceil+2.$
Theorem 2.1.
Let $f$ be as in (1.2) with $a_{1},\ldots,a_{r}\in{\mathbb{F}}_{p}^{*}$. Assume that there exists $\varepsilon$ with $03/92\geq\varepsilon>$ 0 and such that for all $1\leq i\neq j\leq r$:
$\gcd(e_{i},p-1)\leq 0.5p^{7/26+14\varepsilon/13},$
$\gcd(e_{i}-e_{j},p-1)\leq p^{8/13+32\varepsilon/13},$
$\gcd(e_{i},e_{j},p-1)\leq p^{3/26-46\varepsilon/13}.$
Then, we have
$S_{p}(f)\ll p^{1-\kappa_{r}(\varepsilon)}.$
Remark 2.2.
There could be several variants of Theorem 2.1 with slightly relaxed conditions on the corresponding greatest common divisors but with weaker estimates. Since, generically, the greatest common divisor of two integers tends to be small, we imposed the above more stringent conditions, which however maximise the strength of the bound.
Furthermore, we also have a simpler version of Theorem 2.3 with $\varepsilon=3/92$, restricting to $\gcd(e_{i},p-1)=1$, while relaxing the condition on $\gcd(e_{i}-e_{j},p-1)$.
It is convenient to define
(2.4) $\rho_{r}=\kappa_{r}(3/92).$
Theorem 2.3.
Let $f$ be as in (1.2) with $a_{1},\ldots,a_{r}\in{\mathbb{F}}_{p}^{*}$. Assume that for all $1\leq i\neq j\leq r$:
$\gcd(e_{i},p-1)=1\qquad\mbox{and}\qquad\gcd(e_{i}-e_{j},p-1)\leq p^{16/23}.$
Then, we have
$S_{p}(f)\ll p^{1-\rho_{r}}.$
Remark 2.4.
Using Theorems 2.1 and 2.3 combined with the Koksma–Szüsz inequality, see Lemma 5.1, one can get explicit versions of a result of Bombieri, Bourgain and Konyagin [2, Theorem 17], see also [5, Corollary 9], counting the number of solutions to the system of congruences
$a_{i}x^{e_{i}}\equiv y_{i}\pmod{p},\qquad x\in{\mathbb{F}}_{p},\ y_{i}\in[A_{i}+1,A_{i}+H_{i}],\ i=1,\ldots,r,$
with some integers $A_{i}$ and $1\leq H_{i}\leq p$.
2.2. Exponential sums with sparse polynomials over ${\mathbb{Z}}_{q}$
We merge the bound of Theorem 2.3 with the argument Bourgain [4] to derive the following explicit version of (1.9) .
We now define
(2.5) $\sigma_{r}=\min\{\kappa_{r}(3/184),1/(50r^{3})\}.$
In fact $\sigma_{r}<\kappa_{r}(3/184)$ only for $r=2$, that is, alternatively we can define
$\sigma_{r}=\begin{cases}\kappa_{r}(3/184),&\text{for}\ r=1\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ r\geq 3,\\ 1/400,&\text{for}\ r=2.\end{cases}$
Theorem 2.5.
Let $q\geq 2$ be an integer and let $f$ be as  (1.2) and (1.3) of degree $d$. Then uniformly over $d$, we have
$S_{q}(f)\ll\exp\left(d^{18}\right)q^{1-\sigma_{r}+o(1)},\qquad\text{as}\ q\to\infty,$
where $\sigma_{r}$ is given by (2.5).
Remark 2.6.
We emphasise that the main point of Theorem 2.5 is that the degree $d$ may grow with $q$, as the saving $\sigma_{r}$ and the implied constant depend only on the sparsity $r$ rather than on $d$, while the rate of decay of $o(1)$ in the exponent of $q$ depends only on $q$. Clearly, for a fixed $d$, the bound (1.8) of [36, Theorem 1] is much stronger.
Remark 2.7.
The power $18$ of $d$ in Theorem 2.5 is not optimised and can easily be reduced and the cost of redefining and reducing $\sigma_{r}$ as well. We have also ignored the factor $1/50$ at the front of $d^{18}$ which our argument gives, see the final estimate in the proof of Theorem 2.5.
Remark 2.8.
The limit of our technique with respect to the dependence on $d$ is $d^{26/3+\varepsilon}$, with any $\varepsilon>0$, instead of $d^{18}$ in Theorem 2.5 due to (4.3c). However, if we impose that the exponents $e_{i}$ are pairwise coprime, this issue disappears, leaving (4.3a) as the bottleneck. Thus, tightening (4.3a) and (4.3b) according to the conditions in Theorem 2.1, provides a much smaller exponent than $26/3$.
2.3. Multidimensional distribution of the power generator
We recall that the discrepancy $D(\Gamma)$ of a sequence $\Gamma=(\bm{\gamma}_{n})_{n=1}^{N}$ of $N$ points in the $s$-dimensional unit cube $[0,1]^{s}$, is defined by
$D(\Gamma)=\sup_{{\mathcal{B}}\subseteq[0,1]^{s}}\left|\frac{N({\mathcal{B}})}{N}-(\beta_{1}-\alpha_{1})\cdots(\beta_{s}-\alpha_{s})\right|,$
with the supremum is taken over all boxes
${\mathcal{B}}=[\alpha_{1},\beta_{1}]\times\ldots\times[\alpha_{s},\beta_{s}]\subseteq[0,1)^{s},$
and where $N({\mathcal{B}})$ is the number of elements of $\Gamma$, which fall in ${\mathcal{B}}$, and
${\mathrm{vol\,}}{\mathcal{B}}=(\beta_{1}-\alpha_{1})\cdots(\beta_{s}-\alpha_{s})$
is the volume of ${\mathcal{B}}$.
Furthermore, for an integer $\vartheta\not\equiv 0\pmod{p}$, we denote by $D_{e,\vartheta,s}(p,N)$ the discrepancy of the points
$\left(\frac{u_{n}}{p},\ldots,\frac{u_{n+s-1}}{p}\right),\qquad n=1,\ldots,N,$
generated by (1.10) with $u_{0}=\vartheta$.
For $s=1$, a bound on $D_{e,\vartheta,1}(p,N)$ is given by [27, Corollary 3]. However, the approach of [27] does not extend to $s\geq 2$. Here we use Theorem 2.3 to establish such a result.
Clearly, if $\gcd(e,p-1)=1$ and $\gcd(\vartheta,p)=1$, then it is easy to see that the sequence (1.10) is purely periodic.
Theorem 2.9.
Let $\gcd(\vartheta,p)=1$ be of multiplicative order $T$ modulo $p$ and let $\tau$ be the period of the sequence (1.10). Then, for any fixed $s\geq 1$, for $N\leq\tau$ we have the following estimate
$D_{e,\vartheta,s}(p,N)\leq N^{-1/2}p^{1/2-0.5\rho_{2s}+o(1)},$
where $\rho_{r}$ is given by (2.4).
Obviously, the bound of Theorem 2.9 is nontrivial for $\tau\geq N\geq p^{1-\rho_{2s}+\varepsilon}$ with any fixed $\varepsilon>0$.
Remark 2.10.
It is easy to see that the bound of Theorem 2.9 also holds for the discrepancy of the points
$\left(\frac{u_{n+h_{1}}}{p},\ldots,\frac{u_{n+h_{s}}}{p}\right),\qquad n=1,\ldots,N,$
corresponding to arbitrary shifts $0\leq h_{1}<\ldots<h_{s}<\tau$ satisfying $h_{s}-h_{1}=p^{o(1)}$ as $p\to\infty$.
3. Proofs of Theorems 2.1 and 2.3
3.1. Preliminaries
We use the inductive idea of Bourgain [3], however we apply it in the form outlined by Garaev [18, Section 4.4], also applied in [1, 29]. It is important to observe that the basis of induction is formed by the cases $r=1$ and $r=2$ (just $r=1$ is not enough).
We prove both Theorems 2.1 and 2.3 simultaneously, giving more details for the proof of Theorem 2.1, which is more technically involved.
Let us consider the exponential sum $S_{p}(f)$ as in (1.1) with polynomials $f$ as in (1.2). We also use $I_{r,t}$ to denote the number of solutions to the system of equations
(3.1) $x_{1}^{e_{i}}+\cdots+x_{t}^{e_{i}}=y_{1}^{e_{i}}+\cdots+y_{t}^{e_{i}},\qquad i=1,\ldots,r,$
in $x_{1},y_{1},\ldots,x_{t},y_{t}\in{\mathbb{F}}_{p}^{*}$.
Lemma 3.1.
For any integer $s,t\geq 1$ we have
$S_{p}(f)^{2st}\ll p^{2st-2t-2s+r}I_{r,s}I_{r,t}.$
Proof.
For any integer $t\geq 1$, we write
$S_{p}(f)^{t}$ $=\sum_{x_{1},\ldots,x_{t}\in{\mathbb{F}}_{p}}{\mathbf{\,e}}_{p}\left(f\left(x_{1}\right)+\ldots+f\left(x_{t}\right)\right)$
$=\frac{1}{p-1}\sum_{z\in{\mathbb{F}}_{p}^{*}}\sum_{x_{1},\ldots,x_{t}\in{\mathbb{F}}_{p}}{\mathbf{\,e}}_{p}\left(f\left(x_{1}z\right)+\ldots+f\left(x_{t}z\right)\right)$
$=\frac{1}{p-1}\sum_{x_{1},\ldots,x_{t}\in{\mathbb{F}}_{p}}\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\sum_{i=1}^{r}a_{i}z^{e_{i}}\left(x_{1}^{e_{i}}+\cdots+x_{t}^{e_{i}}\right)\right).$
Denoting by $J_{r,t}(\lambda_{1},\cdots,\lambda_{r})$ the number of solutions to the system of equations
$a_{i}\left(x_{1}^{e_{i}}+\cdots+x_{t}^{e_{i}}\right)=\lambda_{i},\qquad i=1,\ldots,r,$
in $x_{1},\ldots,x_{t}\in{\mathbb{F}}_{p}^{*}$. we derive
$|S_{p}(f)|^{t}$ $\ll p^{-1}\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\cdots,\lambda_{r})\left|\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\lambda_{1}z^{e_{1}}+\cdots+\lambda_{r}z^{e_{r}}\right)\right|.$
Since $a_{1},\ldots,a_{r}\neq 0$, we have
$\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\ldots,\lambda_{r})=(p-1)^{t},$
$\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\ldots,\lambda_{r})^{2}=I_{r,t}.$
Next, writing
$J_{r,t}(\lambda_{1},\ldots,\lambda_{r})=J_{r,t}(\lambda_{1},\ldots,\lambda_{r})^{(s-1)/s}\left(J_{r,t}(\lambda_{1},\ldots,\lambda_{r})^{2}\right)^{1/(2s)}$
and applying the Hölder inequality, we derive
$\left|S_{p}(f)\right|^{2st}$ $\ll p^{-2s}\left(\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\cdots,\lambda_{r})\right)^{2s-2}$
$\qquad\ \qquad\qquad\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}J_{r,t}(\lambda_{1},\ldots,\lambda_{r})^{2}$
$\qquad\qquad\qquad\qquad\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}\left|\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\lambda_{1}z^{e_{1}}+\cdots+\lambda_{r}z^{e_{r}}\right)\right|^{2s}$
$\ll p^{2st-2s-2t}I_{r,t}\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}\left|\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\lambda_{1}z^{e_{1}}+\cdots+\lambda_{r}z^{e_{r}}\right)\right|^{2s}.$
Since by the orthogonality of characters we have
$\sum_{\lambda_{1},\ldots,\lambda_{r}\in{\mathbb{F}}_{p}}\left|\sum_{z\in{\mathbb{F}}_{p}^{*}}\mathbf{e}_{p}\left(\lambda_{1}z^{e_{1}}+\cdots+\lambda_{r}z^{e_{r}}\right)\right|^{2s}=p^{r}I_{r,s},$
[…]
To apply Lemma 3.1 inductively, we use the trivial inequality $I_{r,t}\leq I_{r-1,t}$, where $I_{\nu,t}$ is defined as $I_{r,t}$ by with respect to only the first $\nu\leq r$ equations in (3.1). Hence, using the orthogonality of exponential functions, we write
(3.2) $\begin{split}I_{r-1,t}&=\frac{1}{p^{r-1}}\sum_{\lambda_{1},\ldots,\lambda_{r-1}\in{\mathbb{F}}_{p}}\left|\sum_{y\in\mathbb{F}_{p}^{*}}\mathbf{e}_{p}({\lambda_{1}y^{e_{1}}+\cdots+\lambda_{r-1}y^{e_{r-1}}})\right|^{2t}\\ &=p^{2t-r+1}+O\left(\Delta\right),\end{split}$
with
$\Delta=\frac{1}{p^{r-1}}\sum_{\begin{subarray}{c}\lambda_{1},\ldots,\lambda_{r-1}\in{\mathbb{F}}_{p}\\ (\lambda_{1},\ldots,\lambda_{r-1})\neq\mathbf{0}\end{subarray}}\left|\sum_{y\in\mathbb{F}_{p}^{*}}\mathbf{e}_{p}({\lambda_{1}y^{e_{1}}+\cdots+\lambda_{r-1}y^{e_{r-1}}})\right|^{2t},$
where $\mathbf{0}$ is a zero vector in ${\mathbb{F}}_{p}^{r-1}$. Thus we estimate $I_{r,t}$ via exponential sums with $(r-1)$-sparse polynomials, which enable inductive approach. Furthermore, for $t=2$ we use the following bound which is a weakened form or a result of Cochrane and Pinner [9, Theorem 7.1].
Lemma 3.2.
For any $\varepsilon>0$ we have
$I_{2,2}\ll p^{3-4\varepsilon},$
provided that
$\gcd(e_{1},p-1),\,\gcd(e_{2},p-1)\leq 0.5p^{7/26+14\varepsilon/13},$
$\gcd(e_{1}-e_{2},p-1)\leq p^{8/13+32\varepsilon/13},$
$\gcd(e_{1},e_{2},p-1)\leq p^{3/26-46\varepsilon/13}.$
Proof.
Let $d=\gcd(e_{1},e_{2},p-1)$. Assume that $p$ is sufficiently large. By [9, Theorem 7.1] we have the bound
(3.3) $I_{2,2}\ll d^{26/23}p^{66/23},$
provided that (with slightly relaxed constants)
(3.4) $\begin{split}&\gcd(e_{1},p-1),\gcd(e_{1}-e_{2},p-1)\leq(p/d)^{16/23},\\ &\gcd(e_{2},p-1)\leq 0.5(p/d)^{7/23}.\end{split}$
[…]
Using the same argument as in the proof of Lemma 3.2, while taking $d=1$, we also see that [9, Theorem 7.1] leads to the following simpler version of Lemma 3.2.
Lemma 3.3.
We have
$I_{2,2}\ll p^{3-3/23},$
provided that
$\gcd(e_{1},p-1)=\gcd(e_{2},p-1)=1,$
$\gcd(e_{1}-e_{2},p-1)\leq p^{16/23}.$
3.2. Induction basis
First we observe that the bounds (1.5) and (1.6) ensure that under the condition of Theorems 2.1 and 2.3 the desired results hold for $r=1$ and $r=2$.
Indeed, for $r=1$, under the assumption of $e_{1}$ in Theorem 2.1 the bound (1.5) implies
$S_{p}(f)\ll d^{1/2}p^{2/3}(\log p)^{1/6}\ll\left(p^{7/26+14\varepsilon/13}\right)^{1/2}p^{2/3}(\log p)^{1/6}\ll p^{1-\varepsilon},$
since for $\varepsilon<31/240$ we have
$\frac{1}{2}\left(7/26+14\varepsilon/13\right)+2/3<1-\varepsilon.$
Furthermore, under the assumption of $e_{1}$ in Theorem 2.3 we obviously have $S_{p}(f)=0$. Hence, we can take $\kappa_{1}(\varepsilon)=\varepsilon$.
Next, for $r=2$, under the assumption of $e_{1}$ and $e_{2}$ in Theorem 2.1 the bound (1.6) implies
$S_{p}(f)\ll p^{8/13+32\varepsilon/13}+\left(p^{3/26-46\varepsilon/13}\right)^{13/46}p^{89/92}\ll p^{1-\varepsilon}$
since for $\varepsilon\leq 1/9$ we have
$\max\left\{8/13+32\varepsilon/13,\frac{13}{46}\left(3/26-46\varepsilon/13\right)+89/92\right\}\leq 1-\varepsilon.$
It remains to observe that $\min\{31/240,1/9\}>3/92$.
Finally, under the the assumption of $e_{1}$ and $e_{2}$ in Theorem 2.3, by (1.6), we have
$S_{p}(f)\ll p^{16/23}+p^{89/92}\leq p^{89/92}=p^{1-3/92}.$
Hence, recalling the definition (2.1), we see that this establishes the basis of induction for Theorems 2.1 and 2.3.
3.3. Induction step
We now assume that $r\geq 3$ and that the desired result holds for all $k$-sparse polynomials with $k<r$ and establish it for $r$-sparse polynomials.
Let $\Delta$ be as in (3.2). By the inductive assumption, we have
$\Delta$ $=\sum_{\begin{subarray}{c}\lambda_{1},\ldots,\lambda_{r-1}\in{\mathbb{F}}_{p}\\ (\lambda_{1},\ldots,\lambda_{r-1})\neq\mathbf{0}\end{subarray}}\left|\sum_{y\in\mathbb{F}_{p}^{*}}\mathbf{e}_{p}({\lambda_{1}y^{e_{1}}+\cdots+\lambda_{r-1}y^{e_{r-1}}})\right|^{2t}$
$\leq p^{2(t-2)(1-\kappa_{r-1}(\varepsilon))}\sum_{\begin{subarray}{c}\lambda_{1},\ldots,\lambda_{r-1}\in{\mathbb{F}}_{p}\\ (\lambda_{1},\ldots,\lambda_{r-1})\neq\mathbf{0}\end{subarray}}\left|\sum_{y\in\mathbb{F}_{p}^{*}}\mathbf{e}_{p}({\lambda_{1}y^{e_{1}}+\cdots+\lambda_{r-1}y^{e_{r-1}}})\right|^{4}$
$\leq p^{2(t-2)(1-\kappa_{r-1}(\varepsilon))}I_{r-1,2}\leq p^{2(t-2)(1-\kappa_{r-1}(\varepsilon))}I_{2,2}.$
Hence, by Lemma 3.2 we see that
$\Delta\ll p^{2(t-2)(1-\kappa_{r-1}(\varepsilon))+3-\varepsilon}=p^{2t-1-2(t-2)\kappa_{r-1}(\varepsilon))-\varepsilon}.$
Then we see from (3.2) that
$I_{r-1,t}\ll p^{2t-r+1}+p^{2t-1-2(t-2)\kappa_{r-1}(\varepsilon))-\varepsilon}\ll p^{2t-r+1}$
provided
(3.5) $2(t-2)\kappa_{r-1}(\varepsilon)\geq r-2-\varepsilon.$
Hence, under the condition (3.5), by Lemma 3.1, taken with $s=2$, we have
$S_{p}(f)^{4t}\ll p^{4t-3}\,I_{r-1,2}\leq p^{4t-3}\,I_{2,2}.$
Since under the conditions of Theorem 2.1, the conditions of Lemma 3.2 are also satisfied, we derive
$S_{p}(f)^{4t}\ll p^{4t-4\varepsilon}.$
Taking
$t=t_{r}(\varepsilon),$
where $t_{r}(\varepsilon)$ is given by (2.3), to ensure that (3.5) holds, we derive $S_{p}(f)\ll p^{1-\varepsilon/t_{r}(\varepsilon)}=p^{1-\kappa_{r}(\varepsilon)}$, where $\kappa_{r}(\varepsilon)$ is as in (2.2).
Finally, we see that under the conditions of Theorem 2.3 we can apply Lemma 3.3 and check that the above bounds hold with $\varepsilon=3/92$.
4. Proof of Theorem 2.5
4.1. Exponential sums modulo prime powers
We first recall the following bound on the number of zeros of sparse polynomials in ${\mathbb{F}}_{p}$, which follows instantly from [6, Lemma 7].
Lemma 4.1.
Let $p$ be any prime and let $f\in{\mathbb{F}}_{p}[X]$ be be as in (1.2) and (1.3) with $q=p$. Let
$D=\max_{1\leq i\neq j\leq r}\gcd(e_{i}-e_{j},p-1).$
Then, the equation
$f(X)=0,\qquad x\in{\mathbb{F}}_{p},$
has at most $O\left(p^{1-1/r}D^{1/r}\right)$ solutions.
We now estimate
Lemma 4.2.
Let $p$ be any prime and let $m\geq 2$ be any integer. Assume that $f$ is as in (1.2) satisfying (1.3) with $q=p$. Assume that for all $1\leq i\leq r$ we have
(4.1) $\gcd(e_{i},p^{m})\leq p^{m/3}.$
Then, we have
$S_{p^{m}}(f)\ll p^{m-1/r}D^{1/r}$
where $D$ is as in Lemma 4.1.
Proof.
We start as in the proof of [4, Proposition 8]. Let us first set,
$n=\left\lceil m/2\right\rceil.$
Writing
$S_{p^{m}}(f)=\frac{1}{p^{n}}\sum_{x\in{\mathbb{Z}}_{p^{m}}^{*}}\sum_{1\leq y\leq p^{n}}\mathbf{e}_{p^{m}}(f(x+p^{n}y)),$
and noting that
$f(x+p^{n}y)=f(x)+p^{n}f^{\prime}(x)y\pmod{p^{m}},$
we have
$|S_{p^{m}}(f)|$ $\leq\frac{1}{p^{n}}\sum_{x\in{\mathbb{Z}}_{p^{m}}^{*}}\left|\sum_{1\leq y<p^{n}}\mathbf{e}_{p^{m-n}}(f^{\prime}(x)y)\right|$
=#$=\#\left\{x\in{\mathbb{Z}}_{p^{m}}^{*}:\leavevmode\nobreak\ f^{\prime}(x)\equiv 0\pmod{p^{m-n}}\right\}.$
We now write
$f^{\prime}(x)=\sum_{i=1}^{r}a_{i}e_{i}x^{e_{i}-1}=p^{u}\sum_{i=1}^{r}b_{i}x^{e_{i}-1},$
with some integers $u$ and $b_{1},\ldots,b_{r}$ satisfying $\gcd(b_{1},\ldots,b_{r},p)=1$. Since $m\geq 2$, the condition (4.1) implies that $u\leq m/3<m-n$. In particular,
$f^{\prime}(x)\equiv 0\pmod{p^{m-n}}$
implies
$\sum_{i=1}^{r}b_{i}x^{e_{i}-1}\equiv 0\pmod{p}.$
[…]
4.2. Concluding the proof
If
$q=\prod_{p\mid q}p^{m_{p}}$
is the prime factorisation of $q$, then one can write
(4.2) $S_{q}(f)=\prod_{p\mid q}S_{p^{m_{p}}}(\lambda_{p}f),$
with some integers $\lambda_{p}$ such that $\gcd(\lambda_{p},p)=1$, see, for example, [21, Equation (12.21)].
We consider the following system of three independent conditions:
(4.3a) $\gcd(e_{j},p-1)\leq p^{1/4},\qquad 1\leq u\leq r,$
(4.3b) $\gcd(e_{i}-e_{j},p-1)\leq p^{3/5},\qquad 1\leq i\neq j\leq r,$
(4.3c) $\gcd(e_{i},e_{j},p-1)\leq p^{3/52},\qquad 1\leq i\neq j\leq r.$
Note that the condition (4.3b) implies that we have $D\leq p^{3/5}$ in Lemmas 4.1 and 4.2.
Now, we also set
$E=\prod_{1\leq i<j\leq r}(e_{i}-e_{j})$
and consider the sets of primes
${\mathcal{P}}_{1}=\left\{p:\leavevmode\nobreak\ \gcd\left(p^{m_{p}},e_{1}\ldots e_{r}\right)>p^{m_{p}/3}\right\},$
${\mathcal{P}}_{2}=\left\{p:\leavevmode\nobreak\ m_{p}>20r^{2}\text{ and }E\equiv 0\pmod{p^{\lfloor m_{p}/(10r)\rfloor}}\right\},$
${\mathcal{P}}_{3}=\bigl{\{}p\not\in{\mathcal{P}}_{1}:\leavevmode\nobreak\ m_{p}\leq 20r^{2}\text{ and }p\text{ does not satisfy}$
[…]
Let us define
$q_{i}=\prod_{p\in{\mathcal{P}}_{i}}p^{m_{p}},\qquad i=1,2,3,$
and follow the arguments as in the proof of in [4, Proposition 11].
Note that
$|E|\leq d^{r(r-1)/2}<d^{r^{2}/2}.$
Hence for $q_{1}$ and $q_{2}$ we have the estimates
(4.4) $q_{1}<d^{3r}\qquad\mbox{and}\qquad q_{2}<d^{20r^{3}},$
see [4, Equations (3.11) and (3.12)].
For any $p\in{\mathcal{P}}_{3}$, failing one of the conditions (4.3a), (4.3b) and (4.3c) means
$p^{1/4}\leq d,\quad\text{or}\quad p^{3/5}\leq d,\quad\text{or}\quad p^{3/52}\leq d.$
Hence, in any case we have $p\leq d^{52/3}$. Therefore, we trivially have
(4.5) $q_{3}<\prod_{p\leq d^{52/3}}p^{20r^{2}}.$
By the classical result of Rosser and Schoenfeld [31, Theorem 9], we see that (4.5) implies
(4.6) $q_{3}<\exp\left(1.02\cdot 20r^{2}d^{52/3}\right)\leq\exp\left(21r^{2}d^{52/3}\right)$
Next, taking ${\mathcal{P}}={\mathcal{P}}_{1}\cup{\mathcal{P}}_{2}\cup{\mathcal{P}}_{3}$, it follows from (4.2) that
(4.7) $S_{q}(f)\ll q_{1}q_{2}q_{3}\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}}\\ p\mid q\end{subarray}}\left(\sum_{x\in{\mathbb{Z}}_{p^{m_{p}}}^{*}}\mathbf{e}_{p^{m_{p}}}(\lambda_{p}f(x))\right).$
We now follow the argument as in the proof of [4, Proposition 11], and decompose the product of the sums on the right above as the product of three sums according to,
$m_{p}=1,\qquad 1<m_{p}\leq 20r^{2},\qquad m_{p}>20r^{2},$
and denote these products $S_{1},S_{2}$, and $S_{3}$ respectively
To estimate $S_{1}$, instead of the conditions of Theorem 2.1 corresponding to $\varepsilon=3/184$ we impose more stringent conditions (4.3a), (4.3b) and (4.3c), which guarantee that the bound of Theorem 2.1 holds with $\kappa_{r}(3/184)\geq\sigma_{r}$, as defined in (2.5), and we derive
(4.8) $S_{1}\ll Q_{1}^{1-\sigma_{r}+o(1)},\quad\mathrm{where}\ Q_{1}=\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}},\leavevmode\nobreak\ p\mid q\\ m_{p}=1\end{subarray}}p,$
where, here and below, the term $Q^{o(1)}$ from the classical bound on the divisor function [21, Equation (1.81)].
Moreover, applying Lemma 4.2 for $m_{p}\geq 2$, Since for $p\not\in{\mathcal{P}}$ with $m_{p}\geq 2$, the condition (4.3b) that we have $D\leq p^{3/5}$. Hence
(4.9) $S_{2}\ll Q_{2}^{1-1/(50r^{3})+o(1)},\quad\mathrm{where}\ Q_{2}=\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}},\leavevmode\nobreak\ p\mid q\\ 2\leq m_{p}\leq 20r^{2}\end{subarray}}p^{m_{p}}.$
Finally to estimate $S_{3}$, since we run over $m_{p}>20r^{2}$ while $p\not\in{\mathcal{P}}_{2}$, we apply [4, Proposition 7], and get
(4.10) $|S_{3}|\leq Q_{3}^{1-1/(10r^{2})+o(1)},\quad\mathrm{where}\ Q_{3}=\prod_{\begin{subarray}{c}p\not\in{\mathcal{P}},\leavevmode\nobreak\ p\mid q\\ m_{p}>20r^{2}\end{subarray}}p^{m_{p}}.$
For any integer $r\geq 1$ we have
$\sigma_{r}\leq\frac{1}{50r^{3}}.$
Combining (4.7), (4.8), (4.9), and (4.10), we have
$S_{q}(f)\ll q_{1}q_{2}q_{3}\left(Q_{1}Q_{2}Q_{3}\right)^{1-\sigma_{r}+o(1)}=q^{1-\sigma_{r}+o(1)}\left(q_{1}q_{2}q_{3}\right)^{\sigma_{r}}.$
Combining (4.4) and (4.6), we see that
$q_{1}q_{2}q_{3}\leq d^{3r+20r^{3}}\exp\left(21r^{2}d^{52/3}\right)\leq\exp\left(r^{3}d^{18}\right),$
concluding the proof.
5. Proof of Theorem 2.9
5.1. Discrepancy and exponential sums
Let $D(\Gamma)$ denote the discrepancy of a sequence $\Gamma=(\bm{\gamma}_{n})_{n=0}^{N}$ of $N$ points in an $s$-dimensional unit cube $[0,1]^{s}$ as defined in Section 2.3.
One of the basic tools used to study uniformity of distribution is the celebrated Koksma–Szüsz inequality [22, 40] (see also [11, Theorem 1.21]), which generalises the Erdős–Turán inequality to arbitrary dimensions. Like the Erdős–Turán inequality, the Koksma–Szüsz inequality gives a bound on the discrepancy via certain exponential sums.
Lemma 5.1.
For any integer $A\geq 1$, we have
$D(\Gamma)\ll\frac{1}{A}+\frac{1}{N}\sum_{\begin{subarray}{c}\mathbf{a}\in\mathbb{Z}^{s}\\ 0<\|\mathbf{a}\|\leq A\end{subarray}}\frac{1}{r(\mathbf{a})}\left|\sum_{n=1}^{N}\mathbf{e}{(\langle\mathbf{a},\bm{\gamma}_{n}\rangle)}\right|,$
where $\langle\mathbf{a},\mathbf{b}\rangle$ denotes the inner product in ${\mathbb{R}}^{s}$,
$\|\mathbf{a}\|=\max_{j=1,\ldots,s}|a_{j}|,\qquad r(\mathbf{a})=\prod_{j=1}^{s}\left(|a_{j}|+1\right),\qquad\mathbf{e}(z)=\exp\left(2\pi iz\right).$
5.2. Poisson summation
We formulate the following result, which is well known and with various modifications has been used in a large number of works. However, since finding a precise general reference is not easy, we present it here with a full proof, which follows closely the argument of Friedlander and Iwaniec [14, Section 3].
Let, as before $\mathbf{e}(z)=\exp\left(2\pi iz\right)$.
Lemma 5.2.
Let $x_{n}$, $n\in{\mathbb{Z}}$ be a both-sided infinite sequence of complex numbers, indexed by integers. Then for any integers $N\geq K\geq 1$, there is a real $\alpha$, such that
$\sum_{n=1}^{N}x_{n}\ll\frac{\log N}{K}\sum_{n=-N}^{N}\left|\sum_{k=1}^{K}x_{n+k}\mathbf{e}(\alpha k)\right|.$
Proof.
As in [14, Section 3], consider the function
$f(x)=\min\{x,1,N+1-x\}$
in the interval $x\in[0,N+1]$ and set $f(x)=0$ elsewhere (that is, the graph of $f(x)$ looks like a trapezoid). Since ${\mathbb{Z}}$ is invariant with respect to shifting by $k\in{\mathbb{Z}}$, we see that
$\sum_{n=1}^{N}x_{n}=\sum_{n\in{\mathbb{Z}}}f(n)x_{n}=\sum_{n\in{\mathbb{Z}}}f(k+n)x_{k+n}.$
Hence, summing over $k=1,\ldots,K$ and changing the order of summation, yields
$\sum_{n=1}^{N}x_{n}=\frac{1}{K}\sum_{n\in{\mathbb{Z}}}\sum_{k=1}^{K}f(k+n)x_{k+n}=\frac{1}{K}\sum_{n=-N}^{N}\sum_{k=1}^{K}f(k+n)x_{k+n}.$
Let
[…]
be the Fourier transform of $f(x)$. Applying Poisson summation to the sum over $n$ gives
$\sum_{n=1}^{N}x_{n}=\frac{1}{K}\int_{-\infty}^{\infty}g(y)\sum_{n=-N}^{N}\mathbf{e}(ny)\sum_{k=1}^{K}e(ky)x_{k+n}dy.$
Therefore
$\sum_{n=1}^{N}x_{n}$ $\ll\frac{1}{K}\int_{-\infty}^{\infty}|g(y)|\sum_{n=-N}^{N}\left|\sum_{k=1}^{K}\mathbf{e}(ky)x_{k+n}\right|dy$
[…]
for some real $\alpha$ for which the maximum is above double sum is achieved. Note that this sum is periodic with period 1 and a continuous function of $y$, so such $\alpha$. We now recall the bound
$\int_{-\infty}^{\infty}|g(y)|\ll\log N,$
[…]
5.3. Multiplicative orders modulo divisors
There is no doubt that the following elementary statement is well-known, it is also given in [27, Lemma 6].
Lemma 5.3.
Let $e$ be of multiplicative order $\tau_{q}$ modulo a positive integer $q$ with $\gcd(e,q)=1$ and of multiplicative order $\tau_{r}$ modulo a positive divisor $r\mid q$. Then
$\tau_{r}\geq\frac{r}{q}\tau_{q}.$
The following lower bound follows immediately from Lemma 5.3, and is also given as [27, Lemma 7].
Lemma 5.4.
Let $e$ be of multiplicative order $\tau_{q}$ modulo a positive integer $q$ with $\gcd(e,q)=1$. Then for any positive integer $h$,
$\gcd(e^{h}-1,q)\leq hq/\tau_{q}.$
Next, we apply Lemma 5.3 to derive the following variant of [27, Lemma 7].
Lemma 5.5.
Let $e$ be of multiplicative order $\tau_{q}$ modulo a positive integer $q$ with $\gcd(e,q)=1$ For a fixed integer $m\geq 0$ there at most $qK/(r\tau_{q})+1$ positive integers $k\leq K$, with
$\gcd(e^{k}-e^{m},q)=r.$
Proof.
[…]
5.4. Bounding exponential sums
In order to use Lemma 5.1 we need to estimate the exponential sums
$S_{\mathbf{a}}(N)=\sum_{n=1}^{N}{\mathbf{\,e}}_{p}\left(a_{1}u_{n}+\cdots+a_{s}u_{n+s-1}\right),$
with $\mathbf{a}=(a_{1},\ldots,a_{s})\in{\mathbb{F}}_{p}^{s}$, for which we prove that
(5.1) $|S_{\mathbf{a}}(N)|\ll\left(\sqrt{NT/\tau}p^{7/46}+N^{1/2}p^{1/2-0.5\rho_{2s}}\right)p^{o(1)}$
provided that $\mathbf{a}$ is a non-zero vector.
We see from (1.10) that
$u_{n}\equiv\vartheta^{e^{n}}\pmod{p}.$
Since $\gcd(e,T)=1$, this allows us to extend the definition of $u_{n}$ for all $n\in{\mathbb{Z}}$.
By Lemma 5.2, for some real $\alpha$ we have
$S_{\mathbf{a}}(N)\ll\frac{\log p}{K}\sum_{n=-N}^{N}\left|\sum_{k=1}^{K}{\mathbf{\,e}}_{p}\left(a_{1}u_{n+k}+\cdots+a_{s}u_{n+k+s-1}\right)\mathbf{e}\left(\alpha k\right)\right|.$
In fact we simply choose $K=N$ and by the Cauchy inequality, we have
$S_{\mathbf{a}}(N)^{2}$ $\ll\frac{(\log p)^{2}}{N}\sum_{n=-N}^{N}\left|\sum_{k=1}^{N}{\mathbf{\,e}}_{p}\left(a_{1}u_{n+k}+\cdots+a_{s}u_{n+k+s-1}\right)\mathbf{e}\left(\alpha k\right)\right|^{2}$
$=\frac{(\log p)^{2}}{N}\sum_{n=-N}^{N}\left|\sum_{k=1}^{N}{\mathbf{\,e}}_{p}\left(\sum_{j=1}^{s}a_{j}\vartheta^{e^{n+k+j}}\right)\mathbf{e}\left(\alpha k\right)\right|^{2}.$
Since $N\leq\tau$, all elements $\vartheta^{e^{n}}$, $n=1,\ldots,N$, are distinct elements of ${\mathbb{F}}_{p}$. Hence in the sequence $\vartheta^{e^{n}}$, $n=0,\pm 1,\ldots,\pm N$, each element of ${\mathbb{F}}_{p}$ occurs at most $3$ times. Therefore
$S_{\mathbf{a}}(N)^{2}\ll\frac{(\log p)^{2}}{N}\sum_{x\in{\mathbb{F}}_{p}^{*}}\left|\sum_{k=1}^{N}{\mathbf{\,e}}_{p}\left(\sum_{j=1}^{s}a_{j}x^{e^{k+j}}\right)\mathbf{e}\left(\alpha k\right)\right|^{2}$
and after squaring out and opening up, we arrive to
(5.2) $S_{\mathbf{a}}(N)^{2}\ll\frac{(\log p)^{2}}{N}\sum_{k,m=1}^{N}\left|\sum_{x\in{\mathbb{F}}_{p}^{*}}{\mathbf{\,e}}_{p}\left(\sum_{j=1}^{s}a_{j}\left(x^{e^{k+j}}-x^{e^{m+j}}\right)\right)\right|.$
In order to apply Theorem 2.3 we need to control the greatest common divisors for the differences of the exponents of the polynomials in the inner sum in (5.2).
First we observe that $\tau$ is the multiplicative order of $e$ modulo $T$. We can also assume that
$\frac{sT}{\tau}\leq p^{16/23}$
as otherwise the result is trivial. Hence, for $j>i$, by Lemma 5.4
$\gcd\left(e^{k+j}-e^{k+i},p-1\right)$ $=\gcd\left(e^{j-i}-1,T\right)$
$\leq\frac{sT}{\tau}\leq p^{16/23}.$
Next, we fix a divisor $R\mid T$ and observe that by Lemma 5.5 the number of pairs $(k,m)$, $1\leq k,m\leq N$, with
$\max_{i,j=1,\ldots,s}\gcd\left(e^{k+i}-e^{m+j},p-1\right)=R$
is bounded by $O\left(N\left(NT/(R\tau)+1\right)\right)$. Hence, using this for all divisors $R\mid T$ with $R>p^{16/23}$ using the classical bound on the divisor function (see, for example, [21, Equation (1.81)]), we see that the number $U$ of pairs $(k,m)$, $1\leq k,m\leq K$, with
$\max_{i,j=1,\ldots,s}\gcd\left(e^{k+i}-e^{m+j},p-1\right)>p^{16/23}$
can be bounded as
$U\ll N\left(\frac{NT}{p^{16/23}\tau}+1\right)p^{o(1)}.$
For such pairs we estimate the inner sum over $x\in{\mathbb{F}}_{p}$ in (5.2) trivially as $p$ and, since $\mathbf{a}$ is a non-zero vector, we can use the bound of Theorem 2.3 for the remaining pairs.
Thus, we derive
$S_{\mathbf{a}}(N)^{2}$ $\ll\frac{(\log p)^{2}}{N}\left(Up+N^{2}p^{1-\rho_{2s}}\right)$
$\leq\left(\frac{NTp^{7/23}}{\tau}+p+Np^{1-\rho_{2s}}\right)p^{o(1)}.$
Clearly the above bound is trivial if $p^{1-\rho_{2s}}\geq N$, while otherwise
$p\leq Np^{1-\rho_{2s}}.$
Hence the above bound can be simplified as
$|S_{\mathbf{a}}(N)|^{2}\leq\left(\frac{NTp^{7/23}}{\tau}+Np^{1-\rho_{2s}}\right)p^{o(1)},$
and yields (5.1).
In turn, the bound (5.1), together with Lemma 5.1 taken with $A=(p-1)/2$, yields the bound
(5.3) $D_{e,\vartheta,s}(p,N)\leq N^{-1/2}\left(\sqrt{T/\tau}p^{7/46}+p^{1/2-0.5\rho_{2s}}\right)p^{o(1)}.$
Next we observe that the above bound is trivial for $N<p^{1-\rho_{2s}}$. Hence we can assume that $\tau\geq N\geq p^{1-\rho_{2s}}$. In this case
$\sqrt{T/\tau}p^{7/46}\leq\sqrt{p/\tau}p^{7/46}\leq p^{7/46+0.5\rho_{2s}}\leq p^{1/2-0.5\rho_{2s}}$
since $\rho_{2s}\leq\rho_{2}=3/92$.
Thus, the first term in (5.3) never dominates, and the desired result follows.
6. Multidimensional Generalisations
As in [27] one can consider various generalisations of the power generator which arise from the dynamical system generated by iterations of multivariable polynomials.
Let $F_{1},\ldots,F_{m}\in\mathbb{F}_{p}[x_{1},\ldots,x_{m}]$ be $m$ polynomials in $m$ variables over $\mathbb{F}_{p}$. For each $i=1,\ldots,m$ we define the $k$-th iteration of the polynomial $F_{i}$ by the recurrence relation
$F_{i}^{(0)}=X_{i},\quad F_{i}^{(k)}=F_{i}\left(F_{1}^{(k-1)},\ldots,F_{m}^{(k-1)}\right),\quad k=1,2,\ldots.$
More precisely, we define the vectors $u_{n}=(u_{n,1},\ldots,u_{n,m})\in\mathbb{F}_{p}^{m}$ by the recurrence relation
$u_{n+1,i}=F_{i}(u_{n,1},\ldots,u_{n,m}),\quad n=0,1,\ldots,\quad i=1,\ldots,m,$
with some initial vector $\mathbf{=}(u_{0,1},\ldots,u_{0,m})\in\mathbb{F}_{p}^{m}$. The sequence $(\mathbf{u}_{n})$ is eventually periodic with some period $\tau\leq p^{m}$.
Then, we denote $D_{s,\mathbf{F},\mathbf{u}}(p,N)$ the discrepancy of the $N$ points
$\left(\frac{u_{n,1}}{p},\ldots,\frac{u_{n,m}}{p}\ldots,\frac{u_{n+s-1,1}}{p},\ldots,\frac{u_{n+s-1,m}}{p}\right),\quad n=1,\ldots,N,$
inside $[0,1]^{ms}$.
Following [27], we consider the following systems of polynomials.
Type 1
As in [27, Section 3.1], we consider polynomials
$F_{1}=(X_{1}-h_{1})^{e_{1}}G_{1}+h_{1},$
[…]
$F_{m-1}=(X_{m-1}-h_{m-1})^{e_{m-1}}G_{m-1}+h_{m-1},$
$F_{m}=g_{m}(X_{m}-h_{m})^{e_{m}}+h_{m},$
where $G_{i}\in\mathbb{F}_{p}[X_{i+1},\ldots,X_{m}]$ for $i=1,\ldots,m-1$, and $g_{m},h_{i}\in\mathbb{F}_{p}$ for $i=1,\ldots,m$, such that none of the $G_{i}$ has zeroes over $\mathbb{F}_{p}$.
Type 2
The following generalisation of the power generator has been suggested in [27, Section 4]. Let $F_{1}=X_{1}^{e}L_{1}+L_{0}$, where $L_{0},L_{1}$ and the remaining polynomials $F_{2},\ldots,F_{m}$ are all non-constant linear polynomials in ${\mathbb{F}}_{p}[X_{2},\ldots,X_{m}]$.
Type 3
For some integral triangular matrix $\left(e_{i,j}\right)_{i,j=1}^{m}\in\mathbb{Z}^{m\times m}$ of exponents, we consider the system of monomials
$F_{j}=X_{1}^{e_{j,1}}\cdots X_{m}^{e_{j,m}},\qquad j=1,\ldots,m.$
which provides the most natural generalisation of the power generator but also inherits its negative aspects such as homogeneity, see [27, Sections 1.2 and 4].
The argument of [27] allows to get an explicit estimate the discrepancy of the corresponding sequences of vectors only for $s=1$, see, for example, [17, Theorem 3.2] or [27, Corollary 9], and appeal to the result of Bourgain [3, Theorem 1] (that is, to (1.7)), otherwise.
One can use Theorem 2.3 and argue as in [27] to get explicit estimates for the discrepancy $D_{s,\mathbf{F},\mathbf{u}}(p,N)$ for any fixed $s\geq 1$, via the Koksma–Szüsz inequality given in Lemma 5.1.
It is certainly natural to try to extend the above constructions to other systems of polynomials involving more monomials and then use Theorem 2.3. Unfortunately the sparsity of iterations of such polynomials grows exponentially with the number of iterations. So we pose it as an open question to find other polynomial systems to which our new bound of exponential sums applies.
Acknowledgements
The authors would like to thank Arne Winterhof for very helpful comments and suggestions.
During the preparation of this work, the authors were partially supported by the Australian Research Council Grant DP230100530.